# On the Brittleness of CLIP Text Encoders

### v1

@@ Abstract
Multimodal co-embedding models, especially CLIP, have advanced the state of the art in zero-shot classification and multimedia information retrieval in recent years by aligning images and text in a shared representation space. However, such modals trained on a contrastive alignment can lack stability towards small input perturbations. Especially when dealing with manually expressed queries, minor variations in the query can cause large differences in the ranking of the best-matching results. In this paper, we present a systematic analysis of the effect of multiple classes of non-semantic query perturbations in an multimedia information retrieval scenario. We evaluate a diverse set of lexical, syntactic, and semantic perturbations across multiple CLIP variants using the TRECVID Ad-Hoc Video Search queries and the V3C1 video collection. Across models, we find that syntactic and semantic perturbations drive the largest instabilities, while brittleness is concentrated in trivial surface edits such as punctuation and case. Our results highlight robustness as a critical dimension for evaluating vision–language models beyond benchmark accuracy.
@@ 1 Introduction
Our contributions are as follows: First, we present the first large-scale, controlled evaluation of CLIP’s text encoders under structured lexical, syntactic, and semantic perturbations, spanning 190 TRECVID queries and over one million keyframes. Second, we introduce instability as a retrieval-oriented measure of robustness, quantified through overlap@$k$ and RBO, and show that instability is systematic but varies across perturbation classes. Third, we propose a new brittleness index that normalises instability by embedding distance, revealing that CLIP overreacts most strongly to trivial edits (e.g., punctuation, case) rather than to semantic changes. Fourth, by comparing LAION-trained baselines with fine-tuned (FARE2) and alternative (EVA02-L14) text encoders, we show that architectural and training choices improve robustness but do not eliminate brittleness. Finally, we discuss the implications of these findings for real-world applications of CLIP, from retrieval to moderation and safety-critical tasks, and sketch directions for mitigation, including brittleness-aware training objectives and query-normalisation layers.
@@ 4.2 Instability Across Models and Classes
While overlap curves highlight top-$k$ stability, instability provides a graded summary across the full ranking. Figure 2 shows instability distributions aggregated across perturbations: EVA02-L14 and FARE2-H14 are consistently more stable, whereas LAION-based ViTs are more fragile.
@@
Mixed-effects regression confirms these trends quantitatively: (i) baseline instability is substantial ($\hat{\mu}=0.568$, $p\ll 0.001$), (ii) model-level coefficients highlight EVA02-L14 and FARE2-H14 as significantly more stable, and (iii) perturbation-class effects are strongest for Class 2 and Class 3 variants.
@@ 4.3 Brittleness
The brittleness heatmap (Figure 4) highlights a different pattern than raw instability: Class 1 lexical perturbations induce the highest brittleness, since trivial edits (case, punctuation) cause ranking changes disproportionate to their minimal embedding distance. By contrast, larger embedding shifts from Class 2 syntactic or Class 3 semantic perturbations explain much of their instability, yielding lower brittleness. EVA02-L14 and FARE2-H14 consistently exhibit the lowest brittleness, suggesting that training and architecture refinements mitigate, but do not eliminate, overreaction to surface-level noise.

### v2

@@ Abstract
Multimodal co-embedding models, especially CLIP, have advanced the state of the art in zero-shot classification and multimedia information retrieval in recent years by aligning images and text in a shared representation space. However, such modals trained on a contrastive alignment can lack stability towards small input perturbations. Especially when dealing with manually expressed queries, minor variations in the query can cause large differences in the ranking of the best-matching results. In this paper, we present a systematic analysis of the effect of multiple classes of non-semantic query perturbations in an multimedia information retrieval scenario. We evaluate a diverse set of lexical, syntactic, and semantic perturbations across multiple CLIP variants using the TRECVID Ad-Hoc Video Search queries and the V3C1 video collection. Across models, we find that syntactic and semantic perturbations drive the largest instabilities, while brittleness is concentrated in trivial surface edits such as punctuation and case. Our results highlight robustness as a critical dimension for evaluating vision-language models beyond benchmark accuracy.
@@ 1 Introduction
Our contributions are as follows: First, we present the first large-scale, controlled evaluation of CLIP’s text encoders under structured lexical, syntactic, and semantic perturbations, spanning 190 TRECVID queries and over one million keyframes. Second, we introduce instability as a retrieval-oriented measure of robustness, quantified through overlap@$k$ and RBO, and show that instability is systematic but varies across perturbation classes. Third, we propose a new brittleness index that normalises instability by embedding distance, revealing that CLIP overreacts most strongly to trivial edits (e.g., punctuation, case) rather than to semantic changes. Fourth, by comparing LAION-trained baselines with fine-tuned (FARE2) and alternative (EVA02-L14) text encoders, we show that architectural and training choices improve robustness but do not eliminate brittleness. Finally, we discuss the implications of these findings for real-world applications of CLIP, from retrieval to moderation and safety-critical tasks, and sketch directions for mitigation, including brittleness-aware training objectives and query-normalisation layers. The code used for data processing and analysis is available at https://github.com/allie-tran/clip-brittleness.
@@ 4.2 Instability Across Models and Classes
While overlap curves highlight top-$k$ stability, instability provides a graded summary across the full ranking. Figure 2 shows instability distributions aggregated across perturbations: EVA02-L14 and FARE2-H14 are consistently more stable, along with LAION5B, which was trained on multilingual texts.
@@
Mixed-effects regression confirms these trends quantitatively: (i) baseline instability is substantial ($\hat{\mu}=0.568$, $p\ll 0.001$), (ii) model-level coefficients highlight EVA02-L14, FARE2-H14, and LAION5b as significantly more stable, and (iii) perturbation-class effects are strongest for Class 2 and Class 3 variants.
@@ 4.3 Brittleness
The brittleness heatmap (Figure 4) highlights a different pattern than raw instability: Class 1 lexical perturbations induce the highest brittleness, since trivial edits (case, punctuation) cause ranking changes disproportionate to their minimal embedding distance. By contrast, larger embedding shifts from Class 2 syntactic or Class 3 semantic perturbations explain much of their instability, yielding lower brittleness. Depsite having lower instability, LAION5B is remarkably more brittle EVA02-L14 and FARE2-H14 consistently exhibit the lowest brittleness, suggesting that training and architecture refinements mitigate, but do not eliminate, overreaction to surface-level noise.